Reports of 21094
MIF (General)
shinji.miyoki - 9:06 Saturday 03 December 2022 (23117) Print this report
Comment to LO frequency adjustment for AM sideband reduction (23115)

According to the search with the keyword "5.624" or "5.62", we can find the story on how to find and change 5.624*****MHz (= IMC FSR ?). The latest log is klog#10286 in 2019. From 5.6244235 MHz to 5.62443350 MHz, not 5.6246235 whose changing timing is missing.

According to the search with the keyword "IMC FSR", the accuracy of 5.6246235, for example, is much better than the estimated values using IMC. So this value should be finally decided by AM sideband (or offset in the demodulated signals of harmonics) reduction as in this time? While klog#616 shows the comparably precise FSR.

klog#4831 shows how to find the best f1 (3xIMC FSR).

Anyway, IMC mirror positions were actually moved, so the IMC FSR change was expected (We forgot :( ).

IOO (OMC)
kenta.tanaka - 5:10 Saturday 03 December 2022 (23116) Print this report
Beacon Wavefront Sensing with FPMI DARM offset

### Abstract

I implemented DARM offset (~80 pm) to FPMI and locked OMC. OMC transmitted power is ~20mW. Then I excited the drumhead mode of ETMX to generate a 24 kHz beacon signal and tried to observe the beacon in RF17 demodulated signals of OMC REFL WFS. The beacon is found in the demodulated signal. And I made the alignment signal from the beat signal between the beacon and the RF sideband. However, The PIT and YAW signals by beacon WFS seem to have the same shape, and the coherence between them is 1. I'm not sure of the reason, so I need to investigate them.

### What I did

  • I tried to lock OMC during FPMI locked. I implemented some DARM offset in K1:LSC-DARM1_OFFSET. According to klog12763, When PRFPMI was locked, the DARM offset was ~30 pm. So I started the 30 pm first and increased the value gradually. And I tried to lock OMC simultaneously. Finally, I locked OMC when the DARM offset was 80 pm (Fig. 1). The transmitted power was ~20 mW. The power is the same level in O3GK.
    • I wonder if the DARM offset value is lower than I expected. I thought the power at AS from FPMI was 10 times smaller than that from PRFPMI, so I expected the DARM offset for FPMI to become 10 times larger roughly.  But actually, it is not so.
    • In O3GK, we use only one OMCTRANS PD, and there seemed to be some mysterious loss in the path from SR2 to OMC. It may have something to do with that, but I am not sure at this time.
  • After DARM offset implementation and OMC lock. I excited the drumhead mode of ETMX to generate a 24 kHz beacon signal and observed the beacon signal in the RF demodulated signal of OMC REFL WFS. (Fig. 2)
  • I demodulated the 24 kHz signals in RF demodulated signals of each segment of OMC REFL WFS by 24 kHz. (At this time, I implemented the 1mHz cutoff lowpass filters and 24 kHz notch filters in the filter bank of the 24 kHz demodulated for each segment. fig. 3). I made the PIT and YAW signals from the 24 kHz demodulated signal of each segment.
  • Fig. 4 shows the PIT and YAW spectra from Beacon WFS. The peak at 15 Hz is the OMMT2 motion in the YAW direction which I excited intensionally. So I succeeded in the alignment signal from the beat signal 24 kHz between the beacon and the 17 MHz RF sideband for the first time.
  • However, the shape of the spectra of PIT and YAW seems to be the same. I measured their coherence of them. their coherence is 1. So Their signals are strange. I need to do more investigation.

### Next

  • implement DC readout(?) (I already measured the relative gain between OMC and RFPD according to klog12747 (fig. 5))
  • investigate DC centering loop of OMC REFL WFS
  • investigate Beacon WFS signal

 

Images attached to this report
MIF (General)
takafumi.ushiba - 23:04 Friday 02 December 2022 (23115) Print this report
LO frequency adjustment for AM sideband reduction

Abstract:
I changed the frequency of local oscillator for IFO RF sidebands.
RF AM seemed to be reduced by a factor of 10 or more by changing LO frequency from 5.6246235 MHz to 5.62437105 MHz.

Detail:
First, I subtract the dark noise offsets of each PD on POP and REFL for the preparation.
After that, I misaligned both ITM and ITMY for obtaining pure reflection from PRM on REFL PDs.

Then, to reduce the amount of AM sideband as like klog4831klog7281klog9073, and , I modulated the laser frequency.
For the frequency moduration, I used DGCARM path and injected the signal to error point of IMC.
Detailed configurations are as follows:
1. Open switches of SUMMING CMS IN1 and IN2 not to feedback the signal from GR PDH CMS.
2. Change slow output offset of SUMMING CMS to zero.
3. Close CARM CMS IN2 to feedback DGCARM signal to IMC.
4. Close IMC CMS IN2 to add the DGCARM signal to error point of IMC.
5. Turn off all FMs in DGCARM filter bank and injected the excitation signal (1023 Hz, 1000 cnts) from DGCARM filter bank.

Then, I tweeked LO frequency to find the best freqeucy for each PD signal for 1f sideband.
Following table shows the best LO frequencies for each PD.

  Best frequency (MHz)
REFL 45I 5.6243675
REFL 45Q 5.6243728
REFL 56I 5.6243708
REFL 56Q 5.6243731
Average 5.62437105

Since the difference of best frequencies for each PD are larger than the measurement errors (~0.2 Hz), I used the average value this time.
Figure 1 shows spectra of each PD before and after changing LO frequency.
Black and red line show the spectra of each RF PD signal at REFL before and after changing te LO frequency from 5.6246235 MHz to 5.62437105 MHz.
As you can see, AM sideband of 1f decreased by a factor of 10 or more, and AM sideband of 3f became smaller than the PD noise.

After changing the LO, I haven't checked the phase of each PD yet.
So, I will do it for the next step.

Note:
Measured data was stored at /users/Commissioning/data/IFO/2022/1202/.
The LO frequency before changing was 5.6246235 MHz but I cannot find the log to change this frequency except for the miyoki-san's recovery log (klog16155 and klog23010).
So, I don't know from when this value is used.

Images attached to this report
Comments to this report:
shinji.miyoki - 9:06 Saturday 03 December 2022 (23117) Print this report

According to the search with the keyword "5.624" or "5.62", we can find the story on how to find and change 5.624*****MHz (= IMC FSR ?). The latest log is klog#10286 in 2019. From 5.6244235 MHz to 5.62443350 MHz, not 5.6246235 whose changing timing is missing.

According to the search with the keyword "IMC FSR", the accuracy of 5.6246235, for example, is much better than the estimated values using IMC. So this value should be finally decided by AM sideband (or offset in the demodulated signals of harmonics) reduction as in this time? While klog#616 shows the comparably precise FSR.

klog#4831 shows how to find the best f1 (3xIMC FSR).

Anyway, IMC mirror positions were actually moved, so the IMC FSR change was expected (We forgot :( ).

DGS (General)
takahiro.yamamoto - 20:08 Friday 02 December 2022 (23114) Print this report
A new network switch was installed
This work is related to klog#22520.

I installed a new 10G switch at U35 of B1 rack (computer room in the mine) for the test of new data transfer servers.
Writing speed of k1fw1 to hyades-2 didn't improve in this situation.
Thanks to this new network switch, rebooting k1fw1 is not required for switching target disks of hyades-1 and -2.

-----
Originally, k1fw1 and hyades-1 were directly connected.
Because of this, we need to do a cold-boot in order to change a disk server from hyades-1 to hyades-2.
So we was able to do only 1 or 2 tests during one maintenance day.

In order to improve this situation, I installed a new 10G switch at U35 of B1 rack.
(After solving current problem, this switch should be uninstalled due to the rack space, thermal issue, etc.)
k1fw1 and hyades-1 are now connected via the new 10G switch and hyades-2 is also connected to the same switch.

A disk mount is set as follows.
- hyades-1:/data1 => k1fw1:/data1
- hyades-1:/data2 => k1fw1:/data2
- hyades-2:/data => k1fw1:/data9

Original situation is /frames/full, /frames/science, and /frames/trend which are symbolic links to /data2/full, /data2/science /data1/trend, respectively.
In this case, daqd writes down frame files to hyeades-1.

At first, I changed destinations of all three symbolic link as /data9/full, /data9/science, /data9/trend to switch the disk sytem as a new one on hyades-2.
DAQ hang-up was reproced at this time because it takes ~130s for writing 32 second-long data.
So direct connection of k1fw1 and hyades is not a cause of this problem.

As the next step, I set symbolic links as follows.
/frames/full -> /data9/full
/frames/science -> /data2/science
/frames/trend -> /data2/trend
In this case, full frames was written down on hyades-2 and science frames and trend frames were writting down on hyades-1.
DAQ hang-up was reproduced also in this case. But science frames can be written on hyades-1 with ~3s.
A cause of hang-up was the writing time of full frames (~120s).

I also tried to write full frames and science frames to hyades-1 and hyades-2, respectively (swapped destination from above situation).
In this case, writing time on hyades-1 is ~7s (full).
But writing time on hyades-2 is ~110s (science) and daqd hung up.

Because file size of full frames is roughly twice larger than one of science frames.
So writing time on hyades-1 seems to be reasonable (7s vs. 3s).
On the other hand, there seems to be ~100s overhead on hyades-2.

It's better to check the NIC and NFS settings on hyades-2.
VAC (EY)
shoichi.oshino - 18:08 Friday 02 December 2022 (23113) Print this report
Q-mass measurement
DATE : 2022/12/02 08:50:10 UTC
TOTAL PRESSURE : 4.19e-5 Pa

Measured data is stored to /kagra/KAGRA Dropbox/Subsystems/VAC/Q-mass/ETMY/EY_221202_17501.CSV
VAC (EY)
shoichi.oshino - 16:50 Friday 02 December 2022 (23111) Print this report
Network connection of Q-mass at EY
After cabling LAN to the Q-mass(klog 23106), I set up Q-mass at EYC. First I turned dip switch number 7 ON and number 8 OFF. This sets the "fixed IP address" to sensor 2. Then I set an "arbitrary IP address" from QuleeQCS. Finally, I turned on DIP switch number 1 and confirmed that I could connect to Q-mass from the network.
Images attached to this report
VIS (SR2)
fabian.arellano - 16:49 Friday 02 December 2022 (23112) Print this report
Comment to Intercalibration between the seismometer on the ground and the F0 LVDT. (23095)

I added to the plot of the amplitude spectral densities, the SR2 F0 LVDT readout measured for the acceptance check of the Type-B suspensions.

The noise level above 1 Hz was around 150 nm/rtHz and the suspension was in ALIGNED state. This is consistent with the level of 30 nm/rtHz measured recently in TWR_FLOAT state. The excess of noise must be due to injected noise by the control loop.

This is information on the acceptance check measurement:

  • Directory: /users/ryutaro/Log/220802/
  • File: SR2_spe3.xml
  • Date and time: 2022/08/02 at 01:48:48 UTC.
  • Guardian state: ALIGNED (Takahashi-san remembers).
Images attached to this comment
IOO (IMC)
shinji.miyoki - 16:23 Friday 02 December 2022 (23110) Print this report
Comment to Transfer function for IMC servo circuit (23108)

Could you please let us know the actual circuit diagram link?

IOO (IMC)
osamu.miyakawa - 15:48 Friday 02 December 2022 (23108) Print this report
Transfer function for IMC servo circuit

I measured transfer functions of IMC servo circuit since large phase delay was measured by exchanging the wideband EOM.

Measured phase delay at 100kHz was 100 degree order. It sounds too big because of bandwidth of current EOM: 1GHz, and new EOM: 10MHz.

Actually measured phase delay at 100kHz in the servo circuit is less than 10degree, even gain is maximized to +31dB and all of ~10 opamps are used. Attached figure show the actual phase delay with 31dB gain at one amp stage. The opamps are AD829, and they are not so slow.

Actual path has two amp stages, and it means phase can be recovered only up to 20 degree.

Now I suspect a phase flip happened in the new EOM path. Or new EOM has a resonance around 800kHz, so the phase delay could be real. Anyway I will try to add a phase flip next week.

Images attached to this report
Comments to this report:
shinji.miyoki - 16:23 Friday 02 December 2022 (23110) Print this report

Could you please let us know the actual circuit diagram link?

DGS (General)
koji.nakagaki - 14:04 Friday 02 December 2022 (23106) Print this report
LAN wiring at the Yend
[Oshino,Hayakawa,Yoshimura,Yamaguchi,Nakagaki]

We wired LAN cables for Q-Mass and VAC-DAQ at the Yend.
MIF (ITF Control)
takafumi.ushiba - 12:06 Friday 02 December 2022 (23099) Print this report
Comment to PRMI 3F lock trial (23098)

Detail:
Since PRMI 3f lock is not stable, I tried to lock PRMI with carrier on rsonance with 1F signals and checked the 3F signal.
To lock PRMI with a cariier on resonance, I just flipped the sign of PRCL control and can lock without any problem.
At that time, beam power on POP PDA1 was over 2mW (if the calibration factor in FB is correct), while AS power was about 1.1 mW (fig1).

First, I compared 1f signal and 3f signals when passing the resonance (fig2).
Both 1f and 3f signals don't have strange error-signal shape as reported in klog15181.
I also checked if the zig-zag shape can be reproduced when I rotate the phase and it can be.
So, phase seems fine.

After that, I checked the 3f signal when locking PRMI with 1f signals (fig3).
Power on POP was fluctuated maybe due to the oscillation between S-pol and P-pol reported in klog9325, but it doesn't matter at this moment.
REFL51Q signals are almost zero but REFL135I signals have an offset.

Also, I checked the 1f signal when locking PRMI with 3f signals (fig4).
In this case, POP17 signals (MICH 1f signal) have a huge offset: that means MICH is not locked in the AS dark condition.

Currently, I'm not sure the reason of this offset but one possiblity is RFAM, so we need to check it.

Images attached to this comment
MIF (ASC)
hirose.chiaki - 11:52 Friday 02 December 2022 (23101) Print this report
Set sensor of ASC Yarm using REFL WFS. However, coherence is low.

k.tanka, c.hirose

We considered that the Sensing matrix ratio would not change significantly from day to day. (23087
So I aligned the resonator axis to match the center of the QPDs of TMSY, aligned the incident light axis and the resonator axis with ADS, then record offloaded IMMT2 and PR2 with goodoplev and turned off ADS. Sensing matrices were measured in this state.
(Originally, I would have liked to excite at an earlier frequency with ADS on and make the measurement, but due to the ADS of IMMT2 returning to TM and the OLDAMP of PR2 being a complicated path, I turned ADS off.)
After the ADS stabilized, the DC offset was measured, averaged over a 5-minute period. We have installed an offset to the WFS signal.
(Using script /user/Commissioning/script/asc/Yarm/measure_avg_std.py)

We assume feedback control from the WFS signal back to IMMT2 and PR2.
Set sensor : DHARD->IMMT2, DSOFT->PR2.
When I measured each sensor, it seemed that I could see a signal for each degree of freedom.
However, when the PLANT measurement was performed, the coherence was abnormally low.
I measured the spectrum of the WFS and it does not seem to see the movement of the other mirrors.
The possibilities are the following.

  • The sensing matrix uses the 10Hz result of WFS/oplev with 10Hz excitation of the mirror, but does not consider the phase rotation of the gain switch of the oplev.
  • Since the coherence of QPD1 is slightly lower depending on the degrees of freedom, it is necessary to consider which degrees of freedom should be returned. (Simulation results should also be checked)
  • The IMMT2 oplev signal produces exactly the same spectrum results for PIT and YAW.

Sensing matrix : /user/Commissioning/data/ASC/2022/REFL_WFS/sensmat/20221130/center/TF_ASC_Yarm_REFLWFS_OVLEV_{PR2,IMMT2}_{P,Y}.xml 
PLANT : /user/Commissioning/data/ASC/2022/REFL_WFS/PLANT/20221130/PLANT_{IMMT2,PR2}_{DHARD,DSOFT}_{PIT,YAW}.xml
measure sensor   : /user/Commissioning/data/ASC/2022/REFL_WFS/sensmat/20221130/{DHARD,DSOFT}_{P,Y}.xml
spectrum: /user/Commissioning/data/ASC/2022/REFL_WFS/PLANT/20221130/spc_PR2_IMMT2.xml

Images attached to this report
DGS (General)
satoru.ikeda - 11:17 Friday 02 December 2022 (23104) Print this report
Updating Model Files

Updating Model Files

Request from Kenta.Tanaka-san
 Model files for k1omc, k1aso, and k1beacon were updated and restarted.
 DAQ restart was also performed.
 

Non-image files attached to this report
VIS (General)
satoru.ikeda - 11:14 Friday 02 December 2022 (23103) Print this report
Comment to Analog Dewhitening BIO Confirmation (22133)

Analog Dewhitening BIO Confirmation

Overview
To check if Analog Dewhitening BIO can control BIO as it may not work in rare cases.

Steps
 $ cd /opt/rtcds/userapps/release/vis/common/scripts/biocheck/
 $ biocheck.py [OPTIC].
 OPTIC=TypeA or TypeB(etc.ETMX,ITMX,BS,SRM...) or ALL
  ALL checks ETMX, ETMY, ITMX, ITMY.BS, SRM, SR2, SR3 in order.

Confirmation result
ETMX:PI:o GAS:o BFV:o BFH:o MNH:o IMH:o IMB:o TM:o MNV:o IMV:o 
ETMY:PI:o GAS:o BFV:o BFH:o MNH:o IMH:o IMB:o TM:o MNV:o IMV:o 
ITMX:PI:o GAS:o BFV:o BFH:o MNH:o IMH:o IMB:x TM:o MNV:o IMV:o 
ITMY:PI:o GAS:o BFV:o BFH:o MNH:o IMH:o IMB:o TM:o MNV:o IMV:o 
BS:TM:o IMV:o PI:o IMH:o GAS:o 
SRM:TM:o IMV:o PI:o IMH:o GAS:o 
SR2:TM:o IMV:o PI:o IMH:o GAS:o 
SR3:TM:o IMV:o PI:o IMH:o GAS:o 

ITMX
 IMB: No change from ON side (OFF in MDEM) (known)
 

Images attached to this comment
Non-image files attached to this comment
VIS (General)
satoru.ikeda - 11:13 Friday 02 December 2022 (23102) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

DGS Regular maintenance day(12/2)

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO
  Check if the script can be started
  Checked that the limit switches installed are green.
 [Check result]
  No problem

Release Sensor
 [Checked the Release Sensor.]
 Checked that Release Sensor is 0, 1: Seated.
 [Check result]
 ETMX, ETMY, ITMX, ITMY : No problem

Picomotor
 [Confirmation]
  PingTest is performed.
 [Result]
  Only known

 NG:
  SR2(IM) : Not restored today due to other work priorities

  MCI,STM1,
  POP_POM,POS_POM,POM1
  MCF(MCO,MCI)
   => Confirmation only, no recovery operation was performed.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF),SR3(IM,BF,STM)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCE,IMMT1,IMMT2,OMMT1,OMMT2,OSTM
  REFL_WFS,POP,POP2,POS,POS2,POP_WFS,AS_WFS
  PCAL(EX1,EX2,EY1,EY2):Not checked because it is ON only when used

Temperature controller
 [Checked]
 Is the temperature acquisition working properly?

 [Check result] 
  Is the temperature acquisition working properly?
  => No problem

Precision air Processer
 [Check Contents]
 Is the temperature acquisition working properly?

 [Result]
  No problem

FunctionGenerator
 (Confirmation)
 Is there a ping response?
 Are all EPICS channels alive?
  [sitemap]-[Commissioning]-[ALS FG]-[Xarm],[Yarm].
 [Check result]
 X and Y PLLs are working properly.

HWP control PC
 (Confirmation)
 Check if it is possible to login to the control PC.
 (Result) 
 Both k1hwp0, 1 and 2 are OK.
 

Images attached to this comment
Non-image files attached to this comment
MIF (ASC)
hirose.chiaki - 10:40 Friday 02 December 2022 (23087) Print this report
Comment to Measurement of stability to sensing matrix of Yarm (23026)

I summarized the results of the sensing matrix.
We did measurement of Sensing matrix in 22, 23, 24, and 28 Nov.
Shift ±4 mm of PIT, YAW direction on ITMY beam spot (add to the set point: PIT's IY, EY by 2 urad: YAW's IY, EY by ±2 urad) from the determined center.
Good oplev of center : following the figure in klog23026.
Good oplev of center on 28th is changing from this work(klog23069), It is possible that the center from the previous day and the center on the 28th may be misaligned.
Replacement of signal generators(klog23008) has reduced the value of the WFS signal.

22th : center, PIT+4mm : IMMT2, PR2
23th : center : IMMT2,PR2,EY
24th : center, YAW±4mm : IY, EY, PR2, IMMT2 (also PR2 and IMMT2 on AS_WFS)
28th : Yaw ±4mm : PR2, IMMT2
We measured the degrees of freedom of each of these three times.
We originally wanted to measure three times on different dates,
but the number of measurements is not enough due to the high volume of measurements and the time it takes for the ADS to stabilize.
Also, PIT-4mm has not yet been measured.

  • The three data fluctuations are large relative to the error calculated from the coherence.
  • The ratio of the two QPD results has not changed significantly in the day-to-day measurements.
    In case the sensing matrices are fluctuating, I will set a larger phase margin for the ASC filter.

Sensing matrix : /user/Commissioning/data/ASC/2022/REFL_WFS/202211{22,23,24,28}/{center,p4mm_PIT, m_4mm_PIT, p4mm_YAW, m4mm_PIT}/TF_ASC_Yarm_REFLWFS_OVLEV_{mirrors}_{P,Y}{,_2,_3}.xml ->I haven't been able to organize some of the files.

Non-image files attached to this comment
MIF (ITF Control)
takafumi.ushiba - 22:45 Thursday 01 December 2022 (23098) Print this report
PRMI 3F lock trial

Continued work from klog23086.

Abstract:
Today, I locked the PRMI with 1f signals with carrier on resonance.
This can be done just flipping the sign of PRCL feedback gain without any problem.
After that, I compared the 1f and 3f signals when PRMI was locked with 3f and 1f signal.
When PRMI was locked by 3f signals, POP17Q (MICH 1f signal) has an offset while REFL135I has an offset when PRMI was locked with 1f signals.
Further investigation is necessary.

Since I have no time to write the detail now, I will post it later.

Comments to this report:
takafumi.ushiba - 12:06 Friday 02 December 2022 (23099) Print this report

Detail:
Since PRMI 3f lock is not stable, I tried to lock PRMI with carrier on rsonance with 1F signals and checked the 3F signal.
To lock PRMI with a cariier on resonance, I just flipped the sign of PRCL control and can lock without any problem.
At that time, beam power on POP PDA1 was over 2mW (if the calibration factor in FB is correct), while AS power was about 1.1 mW (fig1).

First, I compared 1f signal and 3f signals when passing the resonance (fig2).
Both 1f and 3f signals don't have strange error-signal shape as reported in klog15181.
I also checked if the zig-zag shape can be reproduced when I rotate the phase and it can be.
So, phase seems fine.

After that, I checked the 3f signal when locking PRMI with 1f signals (fig3).
Power on POP was fluctuated maybe due to the oscillation between S-pol and P-pol reported in klog9325, but it doesn't matter at this moment.
REFL51Q signals are almost zero but REFL135I signals have an offset.

Also, I checked the 1f signal when locking PRMI with 3f signals (fig4).
In this case, POP17 signals (MICH 1f signal) have a huge offset: that means MICH is not locked in the AS dark condition.

Currently, I'm not sure the reason of this offset but one possiblity is RFAM, so we need to check it.

Images attached to this comment
MIF (General)
hiromi.yasui - 17:22 Thursday 01 December 2022 (23097) Print this report
NEW MEDM SCREENS FOR HWP

I made some MEDM screens for half-wave plate.

sitemap -> SYS -> HWP OVERVIEW

If you have some comments, please let me know.

Images attached to this report
VIS (SR3)
fabian.arellano - 16:53 Thursday 01 December 2022 (23093) Print this report
Intercalibration between the seismometer on the ground and the F0 LVDT.

I calculated the intercalibration factor between the seismometer on the ground and the F0 LVDT, following the method described in entry 22998.

Fig. 1 shows the TF from the seismometer to the F0 LVDT, the coherence and the ASD of the readout of both devices. An advantageous feature of the F0 LVDT in SR3, is its high sensitivity, namely around 2 nm/rtHz above 1 Hz. This allows us to acquire information of the three peaks expected in the TF.

Fig. 2 shows the fit of a transfer function mathematical model to the measured data (see entry 22998). The data used was between 0.1 and 1.5 Hz with a coherence higher than 0.95. In the script, I used the seismometer readout in units of velocity and without a sensor correction filter. 

The value of the intercalibration factor, that must multiply the seismometer readout, is a = 1.0838.

This is information on the Jupyter notebook I used:

  • File name: seismometer_F0_lvdt_intercalibration.ipynb
  • Directory: /kagra/Dropbox/Subsystems/VIS/vis_commissioning/sr3/IP_diag/SENS_corre/Vertical/notebooks/
  • Conda environment: vishack.
  • Diaggui file: sr3_seis2f0_tf_ol_221130.xml
Images attached to this report
VIS (SR2)
fabian.arellano - 16:52 Thursday 01 December 2022 (23095) Print this report
Intercalibration between the seismometer on the ground and the F0 LVDT.

I tried to calculate the intercalibration factor between the seismometer on the ground, and the F0 LVDT. However, given the current microseismic conditions, the LVDT does not seem to have enough sensitivity for this calculation.

The figure shows the TF from the seismometer to the F0 LVDT, the coherence and the ASDs of the readouts of both devices. Additionally, the readout of the SR3 F0 LVDT is also shown for reference (see entry 23093). Such a spectrum was measured simultaneously as the ones for SR2.

As can be easily seen in the plots, the coherence is only high within a very narrow frequency band, and only one of the three expected peaks is visible. Therefore, it does not seem possible to follow the procedure outlined in klog 22998, and used as reported in klog 23093 for SR3. The transfer function is supposed to be close to unity at high frequencies, a value that is highlighted with a horizontal red line.

The limitation is the poor sensitivity of the SR2 LVDT. In the case of SR3, it was possible to calculate the intercalibration factor because its LVDT is more sensitive and the three peaks were clearly visible. Namely, above 1 Hz, the sensitivity of the SR3 LVDT is 15 times higher than that of SR2; the first one reaches 2 nm/rtHz, whereas the second one only 30 nm/rtHz. These levels are indicated with horizontal red lines in the figure.

This is information on the Jupyter notebook and diaggui file I used:

  • File name: seismometer_F0_lvdt_intercalibration.ipynb
  • Directory: /kagra/Dropbox/Subsystems/VIS/vis_commissioning/sr2/IP_diag/SENS_corre/Vertical/notebooks/
  • Conda environment: vishack.
  • Diaggui file: sr2_seis2f0_tf_ol_221130.xml
Images attached to this report
Comments to this report:
fabian.arellano - 16:49 Friday 02 December 2022 (23112) Print this report

I added to the plot of the amplitude spectral densities, the SR2 F0 LVDT readout measured for the acceptance check of the Type-B suspensions.

The noise level above 1 Hz was around 150 nm/rtHz and the suspension was in ALIGNED state. This is consistent with the level of 30 nm/rtHz measured recently in TWR_FLOAT state. The excess of noise must be due to injected noise by the control loop.

This is information on the acceptance check measurement:

  • Directory: /users/ryutaro/Log/220802/
  • File: SR2_spe3.xml
  • Date and time: 2022/08/02 at 01:48:48 UTC.
  • Guardian state: ALIGNED (Takahashi-san remembers).
Images attached to this comment
DetChar (General)
shoichi.oshino - 16:35 Thursday 01 December 2022 (23096) Print this report
Preparation of veto analysis for O4
I start to check the channel list for using veto analysis toward O4.
I checked OpLev channels and modified parameter files of Omicron and Hveto.
These channels were changed after O3GK and also new OpLev (MN, PF) were added.
DetChar (General)
shoichi.oshino - 14:26 Thursday 01 December 2022 (23094) Print this report
LIVE data process was stopped
I found that gw_live_guardian was stopped from 11/25 20:00 UTC.
I resubmit this process to condor and then it starts working.
MIF (General)
takafumi.ushiba - 13:22 Thursday 01 December 2022 (23091) Print this report
Modification of INITIAL_ALIGNMENT guardian

Since GRX and GRY lock are not stable when microseismic is large due to the large feedback signals, I implemented the MCL feedback during he GRX and GRY initial alignment.
Figure 1 and 2 show the feedback signals before and after modification.
Since the liniar range of ALS feedback is between 0.8 and 2.8, so this modification makes the feedback signals small enough.

Images attached to this report
IOO (OMC)
kenta.tanaka - 13:08 Thursday 01 December 2022 (23090) Print this report
Comment to The beacon signal can be observed in OMC REFL WFS RF demod signal but they are too small (23088)

Hirata, Tanaka

We replaced the AI filter with the bypassed AI filter. I changed the channel assigned to Beacon EXC (ch9-12 -> ch 1-4) and IM Damper (ch1-4 -> ch 9-12). (fig .1)

We confirmed the beacon signal in CARM signal became much more significant. (fig. 2). And We confirmed the beacon signal in OMC REFL WFS1 and WFS2 became much more significant. (Just after the confirmation, the earthquake occurred, and OMC died. So I cannot take the photo.)  

### Note 

  • The AC plug for the IO chassis was unplugged accidentally during our work. We replugged it and restarted the IO chassis. And Hirata-san fixed the plug with cable ties (fig .3)
  • The indicator of -18V on the rear panel of the bypassed AI filter doesn't glow (fig. 4). However, the indicators of +/- 15V on the front panel of the filter glow (fig. 5). And ETMX seems to work will. So maybe the LED is broken. 
Images attached to this comment
MIF (General)
satoru.ikeda - 10:41 Thursday 01 December 2022 (23089) Print this report
Comment to Recovery of REFL light power (23051)

Changed log output destination

Old:/kagra/Dropbox/Subsystem/VIS/Scripts/HWP/LogFiles/k1hwp{0,1,2...} _(iSerial).log

New:/kagra/Dropbox/Subsystems/MIF/scripts/HWP/LogFiles/k1hwp{0,1,2...} _(iSerial).log

Search Help
×

Warning

×